# LLM-enhanced Scene Graph Learning for Household Rearrangement

### v1

@@ Abstract.
The household rearrangement task involves spotting misplaced objects in a scene and accommodate them with proper places. It depends both on common-sense knowledge on the objective side and human user preference on the subjective side. In achieving such task, we propose to mine object functionality with user preference alignment directly from the scene itself, without relying on human intervention. To do so, we work with scene graph representation and propose LLM-enhanced scene graph learning which transforms the input scene graph into an affordance-enhanced graph (AEG) with information-enhanced nodes and newly discovered edges (relations). In AEG, the nodes corresponding to the receptacle objects are augmented with context-induced affordance which encodes what kind of carriable objects can be placed on it. New edges are discovered with newly discovered non-local relations. With AEG, we perform task planning for scene rearrangement by detecting misplaced carriables and determining a proper placement for each of them. We test our method by implementing a tiding robot in simulator and perform evaluation on a new benchmark we build. Extensive evaluations demonstrate that our method achieves state-of-the-art performance on misplacement detection and the following rearrangement planning.
journal: TOGcopyright: noneccs: Computer systems organization Robotics
@@ 1. Introduction
Scene analysis is a long-standing problem in computer graphics and vision (Patil et al., 2024) and finds many applications ranging from scene modeling/generation (Fisher et al., 2015) and VR/AR (Li et al., 2021) to various robotic applications (Zhang et al., 2023). Among the applications, some, e.g., robot-operated household rearrangement or room tiding, requires a much more in-depth analysis of the scenes than others (e.g., object goal navigation), due to the involvement of both common-sense knowledge on the objective side and human user preference on the subjective side.
@@
When conducting a room tiding task, the agent needs to spot the misplaced objects and accommodate them in proper places (Sarch et al., 2022). This is quite challenging, especially in an unseen environment, as it relies on a fine-grained and deep understanding of the functionality of the objects. The determination of object functionality often requires a high-order analysis which relies on a larger context (Hu et al., 2018). Moreover, object functionality is often hinged on the house owner’s taste or preference, i.e., the owner’s intent on how an object is being used. For example, the owner could take a cabinet either as a closet or as a bookshelf, depending on where it is located and mainly what objects are placed in.
@@
Large Language Models (LLMs) have been found extremely powerful in challenging, zero-shot task planning including scene rearrangement (Kant et al., 2022; Wu et al., 2023) due to the strong capability of common sense reasoning. However, since the built-in knowledge about object functionality and placement of an LLM is not grounded to the specific scene being considered, scene grounding is usually indispensable to ensure the executability of the generated plans and their relevance to the scene (Rana et al., 2023).
@@
Another important issue to consider is that LLM ignores human preference totally if the prompt is not carefully designed or tuned. In the work of TidyBot (Wu et al., 2023), this is alleviated by explicitly injecting user preference as exemplar placements via prompts and having the LLM to summarize user preference out of the the exemplars. Note that the exemplar placements need to be specific to the scene in consideration to make the summarization scene-grounded, making the exemplar collection tedious. Han et al. (2024) propose an iterative self-training method to align the LLM planner with user preferences, which again relies on user feedback.
@@
In this work, we propose to mine object functionality with user preference alignment directly from the scene itself, without relying on human intervention. With our method, the user instruction can be as simple as “tide the house” and the agent can conduct scene rearrangement automatically based on an in-depth analysis of object functionality and proper reasoning of user preference. One example is shown in Figure 1. Based on our deep analysis of personalized functionality, it appears the user would preferring utilizing the trunk as an unconventional surface for a table. Although this may deviate from the traditional use of a standard trunk, it aligns perfectly with the user’s unique desires for a more tailored and comfortable space.
@@
With AEG, we perform scene rearrangement in two steps. The first step is misplacement detection. For each carriable, we feed its context-induced affordance, its associated receptacle objects and the task instruction into GPT4V. GPT4V then outputs for each carriable a probability of misplacement. In the second step, we choose the top few most probably misplaced carriables and query GPT4V for potential receptacles for correct placement. A straightforward method for receptable query is to put all candidate receptacles in the scene into a very long prompt which may cause degraded accuracy and even hallucinations in output. We, instead, opt to pre-compile task-related affordance (affordance with task description) of all receptacles into an external memory in a similar spirit to Retrieval Augmented Generation (RAG). This way, the prompt for receptacle query is composed of a few most relevant entries retrieved from the memory, thus significantly improving query quality.
Figure 2. Given a scene graph (SG) with key frames, we utilize Large Language Model (LLM) to perform context-induced affordance analysis for all objects in the scene. These affordances are incorporated into the SG, updating the nodes and edges to construct an Affordance Enhanced Graph (AEG). We then evaluate the appropriateness of the current placements of all carriable objects in the AEG based on the affordance information of the objects and their receptacles, identifying misplaced items. For each misplaced carriable, we rate the suitability of each receptacle in the AEG as a placement target. The top k suitable receptacles are selected as candidates, and their affordances are retrieved as prompt to the LLM to generate the placement decision.
We test our method by implementing a robot system for indoor scene rearrangement in the Habitat 3.0 simulator (Puig et al., 2023). Entering an indoor scene, the robot first performs an RGB-D reconstruction and extracts a vanilla scene graph for the scene. It then transforms the scene graph into an AEG and plans actions for the rearrangement task automatically through prompting GPT4V with the scene graph and selected RGB keyframes. We also contribute a benchmark for evaluating the performance of scene rearrangement through annotating the Habitat Synthetic Scenes Dataset (HSSD 200) (Khanna et al., 2023) with ground-truth receptacles for each carriable. Extensive evaluations on existing benchmark and our new benchmark both demonstrate that our method achieves state-of-the-art performance on misplacement detection and the following rearrangement planning in the household rearrangement task.
@@ 2. Related Works
Scene understanding is an important area of research that can be broadly categorized into perception tasks (Georgakis et al., 2017; Silberman et al., 2012; Zhang et al., 2020) and application tasks (Achlioptas et al., 2020; Azuma et al., 2022; Duan et al., 2022). As the field of embodied AI develops, scene understanding research has become more focused on understanding how scenes support the behavior of embodied agents such as human avatars and robots, i.e affordance (Gibson, 1977).
Affordance prediction has been approached at both the object level and the scene level. Object-level methods focuses on understanding how humans use items (Nagarajan et al., 2019; Koppula and Saxena, 2014; Fang et al., 2018; Deng et al., 2021; Nagarajan et al., 2020), while scene-level research emphasizes how to naturally place a human avatar within a scene (Wang et al., 2024; Li et al., 2019; Gupta et al., 2011; Kulal et al., 2023; Grabner et al., 2011; Zhu et al., 2016; Savva et al., 2014, 2016). Different from the aforementioned works which analyze affordance from a human interaction aspect, some studies focus on the analysis object-object affordance (Mo et al., 2022; Tang et al., 2021; LI et al., 2022). Their methods predict the physical interaction between objects, such as placement and stacking. (Tang et al., 2021; LI et al., 2022) extract potential object interactions within a given scene. However, all these object-object affordance works do not consider ¡scene affordance¿ (need to be defined in the introduction) of the objects, which requires an in-depth analysis of factors including the object’s geometry, pose, and the scene context, such as its position in the scene and the objects already interacting with it.
@@ 3. Method
Figure 2 presents a overview of our approach. Starting with a 3D scene, abstracted into a scene graph where each object node is linked to a key frame, our method initiates by harnessing the contextual details of each node. This information is used to augment the graph, infusing it with induced affordances and uncovering new semantic connections between objects. Adopting the objective of ”tidying the house,” the affordance-enriched graph serves as the foundation for identifying objects that are out of place. To determine the most suitable location for each misplaced object, our system initially extracts a set of potential receptacle candidates from the graph. These candidates are then meticulously evaluated and compared to arrive at the optimal placement decision. It is important to highlight that all the pivotal steps of our methodology are executed in a zero-shot learning context, leveraging the advanced capabilities of the Large Language Model (LLM), specifically GPT4V.
Figure 3. Context-induced affordance analysis for objects in the scene graph. (a) For each object, we conduct a local analysis based on its textual contents related to its neighborhood in the graph and its key frame image, storing the results in the nodes as local affordances. (b) We construct object-area-room hierarchies and aggregate the local affordances of all objects within an area. Using LLM, we summarize the content and functionality of the entire area based on the aggregated information, obtaining descriptions for all areas within the room. (c) We use the descriptions of all areas in the room as the global context and assign this context to all nodes within the room to allow the LLM to identify meaningful semantic context information for each object. With this meaningful context, we construct new semantic edges between objects in the same room and update the previously analyzed local affordances, completing the context-induced affordance analysis of the scene graph.
Figure 4. Local affordance analysis of an example receptacle object.
@@ 3.1. Context-induced Affordance Analysis
To simplify the problem setting, we assume that a vanilla scene graph $S$ has already been constructed from a RGBD scanning sequence for the given scene, where each node represents an object instance in the scene and two nodes will be connected with an edge if they are close enough. Moreover, several attributes of the object instance are stored in each node, including the instance ID, category label, located room label, and rearrangement type (carriable, receptacle, or others), where the rearrangement type follows the definitions from the Vocabulary Mobile Manipulation (OVMM)(Yenamandra et al., 2023) problem. In the Household task, only carriable objects can be rearranged, receptacle objects serve as targets for placing carriable objects, and other objects do not interact with the agent during the task but can provide contextual information for other objects. The relationship between two nearby objects is also categories into three different types, including near, on and support, and stored on the connecting edge. Note that if the scene graph is not given, it can be easily constructed following previous works, and details about the graph construction process can be found in the supplementary material.
To help with the inference, we further associate each node with a representative RGB frame selected from the input RGBD scanning sequence to provide a visual context. In more details, for each object instance, we first calculate the number of its projected pixels across all RGB-D images, creating an object-to-image-pixel matrix ($N*M$), where $N$ is the number of object instance and $M$ is the number of frames. To select the most informative frame, we identify all the neighboring objects and sum up their projected pixel numbers together with that of the current instance on each frame. The frame with the maximal number of projected pixels is selected, as it contained most visual information of surrounding context of the instance to be considered.
@@
Local context refers to the neighboring information stored in the scene graph $S$.
For each receptacle and other object in the $S$, we input all the textual contents related to its neighborhood and organize them into a descriptive text with the following template: “The name of this receptacle is [], its category is [], it is located in the room [], its relationships with surrounding objects are: it is near these objects: [], it supports these objects: [], it is supported by these objects: []. ” Note that if the given object does not have a certain type of edge, then the corresponding description will not be generated. Then we combine this template description with its associated key frame, and design a prompt for LLM to analyze the functionality of the object reflected in this scene via Chain of Thought (coT) reasoning. The final output consists of information related to Geometry & Position, Relationships, Functionality, and Fine-grained Category.
Figure 4 shows the local affordance analysis result of an example receptacle. In this example, the LLM first analyzes the geometric features of a table based on the visual information provided by the corresponding image. It combines this information with contextual information given by the scene graph to infer that the table is positioned against the wall at the entrance of the living room. Then, leveraging the position information and the supporting objects of the table in the scene graph, the LLM infers that the table can be likely used for placing outdoor wear. In the functionality module, based on the preceding analysis, the LLM further deduces that the table is suitable for placing items such as shoes that are typically found near entrances, as well as keys and other items that are temporarily stored when entering or leaving. As a conclusion, based on the reasoning above, the LLM assigns a new fine-grained category, ”Entryway organizing console”, to the table based on the above analysis.
Note that when conducting local context analysis, the information enhancement for carriable objects is a bit different, as its local context within the environment frequently changes, and thus we simplify the analysis without consider the position and relationship. For carriable objects, we input the category of object as well as its key frame image, requiring LLM to output the affordance of object with the following information: Geometry & Functionality and Fine-grained Category.
@@
Although our vanilla scene graph $S$ can effectively describe the surrounding context of a receptacle, the functionality of the receptacle may also depends on objects that are far away thus without connecting by an edge in $S$. For example, a sofa and a TV have a strong functional relationship (people sit on the sofa to watch TV), but they are typically arranged at a considerable distance from each other. Therefore, to consider more global context in the scene, we first build a hierarchical structure of $S$ and use it to bridge objects that are spatially distant yet functionally interrelated. In this way, new semantic edges are added, and the information stored in each object node is also updated, resulting in a new affordance-enhanced graph (AEG), as shown in Figure 3.
The details of the hierarchy construction process can be found in the supplementary material. Note that objects are first grouped into an intermediate granularity level called “area”, and then further merged into a room. Using this ”object-area-room” hierarchy to aggregate contextual information is more efficient than aggregating directly from objects to rooms. This method allows for layered, structured storage of aggregated information, preventing the loss of important details.
@@
With the “object-area-room” hierarchy, we first use LLM to aggregate contextual information for each area separately. For each area, we collect the node information stored in $S_{l}$ of all the objects inside the area, and design a prompt for LLM to generate the following content based on a fixed template: Description and Fine-grained name. Therefore, each area will be associated with the above information together with the list of the object inside this area. We then integrate all the information from the areas within a room and use it as the global context for the room to extract distant context for all the receptacle and fixed objects within the room.
@@
For each receptacle, we input the basic spatial information template description, its local affordance analysis as well as the global context of the room to the LLM. With this inputs and a designed prompt, LLM can extract the distant but meaningful context from the global information and add these contexts as semantic edges to the SG. These context output will be in a format of a list of objects, indicating objects within the room that is functionally related to the given receptacle, as well as the specific functional relationship between these objects and the given receptacle. After obtaining this meaningful context, we add semantic directed edges from the given receptacle to related objects, reflecting their functional relations.
@@
In the end, we use LLM to update context-induced affordance based on the previous affordance analysis, the additional global context, as well as the semantic edges, with the same output format as in the local context afforadnce. In this way, our context-induced affordance includes both surrounding spatial context and distant semantic context, resulting in a more comprehensive understanding of receptacle’s functionality in this room. Note that we do not use global context to update the affordance of carriables for the same reason in local analysis. We refer the final scene graph with updated node information and newly added semantic edges as Affordance Enhanced Graph (AEG) and use it for further misplacemnet detection as well as rearrangement planning.
Figure 5. Misplace object scoring with the LLM.

### v2

@@ Abstract.
The household rearrangement task involves spotting misplaced objects in a scene and accommodate them with proper places. It depends both on common-sense knowledge on the objective side and human user preference on the subjective side. In achieving such a task, we propose to mine object functionality with user preference alignment directly from the scene itself, without relying on human intervention. To do so, we work with scene graph representation and propose LLM-enhanced scene graph learning which transforms the input scene graph into an affordance-enhanced graph (AEG) with information-enhanced nodes and newly discovered edges (relations). In AEG, the nodes corresponding to the receptacle objects are augmented with context-induced affordance which encodes what kind of carriable objects can be placed on it. New edges are discovered with newly discovered non-local relations. With AEG, we perform task planning for scene rearrangement by detecting misplaced carriables and determining a proper placement for each of them. We test our method by implementing a tiding robot in simulator and perform evaluation on a new benchmark we build. Extensive evaluations demonstrate that our method achieves state-of-the-art performance in misplacement detection and the following rearrangement planning.
copyright: acmcopyrightccs: Computer systems organization Robotics
@@ 1. Introduction
Scene analysis is a long-standing problem in computer graphics and vision (Patil et al., 2024) and finds many applications ranging from scene modeling/generation (Fisher et al., 2015) and VR/AR (Li et al., 2021) to various robotic applications (Zhang et al., 2023). Among the applications, some, e.g., robot-operated household rearrangement or room tiding, require a much more in-depth analysis of the scenes than others (e.g., object goal navigation), due to the involvement of both common-sense knowledge on the objective side and human user preference on the subjective side.
@@
When conducting a room tiding task, the agent needs to spot the misplaced objects and accommodate them in proper places (Sarch et al., 2022). This is quite challenging, especially in an unseen environment, as it demands a nuanced understanding of each object’s functionality. The determination of object functionality often requires a high-order analysis that considers a broader context (Hu et al., 2018). Moreover, object functionality is often hinged on the house owner’s taste or preference, i.e., the owner’s intent on how an object is being used. For example, the owner could take a cabinet either as a closet or as a bookshelf, depending on where it is located and mainly what objects are placed in. By analyzing the positioning and context of objects, we can infer the homeowner’s preferences and accurately determine the objects’ functionality in real-world indoor scenarios.
@@
Large Language Models (LLMs) have demonstrated remarkable effectiveness in challenging zero-shot task planning tasks, including scene rearrangement (Kant et al., 2022; Wu et al., 2023) due to their strong common-sense reasoning capabilities. However, LLMs inherently lack specific knowledge about object functionality and placement tailored to the scene at hand. Therefore, scene grounding is crucial to ensure the practicality and relevance of the plans they generate (Rana et al., 2023).
@@
Another important issue to consider is that LLM ignores human preference totally if the prompt is not carefully designed or tuned. In the work of TidyBot (Wu et al., 2023), this is alleviated by explicitly injecting user preference as exemplar placements via prompts and having the LLM summarize user preference out of the exemplars. Note that the exemplar placements need to be specific to the scene in consideration to make the summarization scene grounded, making the exemplar collection tedious. Han et al. (2024) propose an iterative self-training method to align the LLM planner with user preferences, which again relies on user feedback. On the other hand, we posit that the existing layout of an indoor scene already encapsulates sufficient human preferences, which can be discerned by analyzing the scene’s context, providing a natural prompt for the agent to perform rearrangement tasks.
@@
In this work, we propose to mine object functionality with user preference alignment directly from the scene itself, without relying on human intervention. With our method, the user instruction can be as simple as “tide the house” and the agent can conduct scene rearrangement automatically based on an in-depth analysis of object functionality and proper reasoning of user preference. One example is shown in Figure 1. Based on our deep analysis of personalized functionality, it appears the user would prefer utilizing the trunk as an unconventional surface for a table. Although this may deviate from the traditional use of a standard trunk, it aligns perfectly with the user’s unique desire for a more tailored and comfortable space.
@@
With AEG, we perform scene rearrangement in two steps. The first step is misplacement detection. For each carriable, we feed its context-induced affordance, its associated receptacle objects, and the task instruction into GPT4V. GPT4V then outputs for each carriable a probability of misplacement. In the second step, we choose the top few most probably misplaced carriables and query GPT4V for potential receptacles for correct placement. A straightforward method for receptacle query is to put all candidate receptacles in the scene into a very long prompt which may cause degraded accuracy and even hallucinations in output. We, instead, opt to pre-compile task-related affordance (affordance with task description) of all receptacles into an external memory in a similar spirit to Retrieval Augmented Generation (RAG). This way, the prompt for receptacle query is composed of a few most relevant entries retrieved from the memory, thus significantly improving query quality.
Figure 2. Given a scene graph (SG) with key frames, we utilize Large Language Model (LLM) to perform context-induced affordance analysis for all objects in the scene. These affordances are incorporated into the SG, updating the nodes and edges to construct an Affordance Enhanced Graph (AEG). We then evaluate the appropriateness of the current placements of all carriable objects in the AEG based on the affordance information of the objects and their receptacles, identifying misplaced items. For each misplaced carriable, we rate the suitability of each receptacle in the AEG as a placement target. The top k suitable receptacles are selected as candidates, and their affordances are retrieved as prompts to the LLM to generate the placement decision.
We test our method by implementing a robot system for indoor scene rearrangement in the Habitat 3.0 simulator (Puig et al., 2023). Entering an indoor scene, the robot first performs an RGB-D reconstruction and extracts a vanilla scene graph for the scene. It then transforms the scene graph into an AEG and plans actions for the rearrangement task automatically by prompting GPT4V with the scene graph and selected RGB keyframes. We also contribute a benchmark for evaluating the performance of scene rearrangement through annotating the Habitat Synthetic Scenes Dataset (HSSD 200) (Khanna et al., 2023) with ground-truth receptacles for each carriable. Extensive evaluations of these two benchmarks demonstrate that our method achieves state-of-the-art performance on misplacement detection and the following rearrangement planning in the household rearrangement task.
@@ 2. Related Works
Scene understanding is an important area of research that can be broadly categorized into perception tasks  (Georgakis et al., 2017; Silberman et al., 2012; Zhang et al., 2020) and application tasks  (Achlioptas et al., 2020; Azuma et al., 2022; Duan et al., 2022). As the field of embodied AI develops, scene understanding research has become more focused on understanding how scenes support the behavior of embodied agents such as human avatars and robots, i.e. affordance  (Gibson, 1977).
Affordance prediction has been approached at both the object level and the scene level. Object-level methods focus on understanding how humans use items  (Nagarajan et al., 2019; Koppula and Saxena, 2014; Fang et al., 2018; Deng et al., 2021; Nagarajan et al., 2020), while scene-level research emphasizes how to naturally place a human avatar within a scene  (Wang et al., 2024; Li et al., 2019; Gupta et al., 2011; Kulal et al., 2023; Grabner et al., 2011; Zhu et al., 2016; Savva et al., 2014, 2016). Different from the aforementioned works which analyze affordance from a human interaction aspect, some studies focus on the analysis of object-object affordance  (Mo et al., 2022; Tang et al., 2021; LI et al., 2022). Their methods predict the physical interaction between objects, such as placement and stacking.  (Tang et al., 2021; LI et al., 2022) extract potential object interactions within a given scene. However, all these object-object affordance works do not consider the ”scene affordance” of the objects, which requires an in-depth analysis of factors including the object’s geometry, pose, and the scene context, such as its position in the scene and the objects already interacting with it.
@@ 3. Method
Figure 2 presents an overview of our approach. Starting with a 3D scene, abstracted into a scene graph where each object node is linked to a key frame, our method initiates by harnessing the contextual details of each node. This information is used to augment the graph, infusing it with induced affordances and uncovering new semantic connections between objects. Adopting the objective of ”tidying the house,” the affordance-enriched graph serves as the foundation for identifying objects that are out of place. To determine the most suitable location for each misplaced object, our system initially extracts a set of potential receptacle candidates from the graph. These candidates are then meticulously evaluated and compared to arrive at the optimal placement decision. It is important to highlight that all the pivotal steps of our methodology are executed in a zero-shot learning context, leveraging the advanced capabilities of the Large Language Model (LLM), specifically GPT4V.
Figure 3. Context-induced affordance analysis for objects in the scene graph. (a) For each object, we conduct a local analysis based on its textual contents related to its neighborhood in the graph and its key frame image, storing the results in the nodes as local affordances. (b) We construct object-area-room hierarchies and aggregate the local affordances of all objects within an area. Using LLM, we summarize the content and functionality of the entire area based on the aggregated information, obtaining descriptions for all the areas. (c) We use the descriptions of all areas in a room as the global context and assign them to all nodes within the room to allow the LLM to identify meaningful semantic context information for each object. With this meaningful context, we construct new semantic edges between objects in the same room and update the affordances, completing the context-induced affordance analysis of the scene graph.
Figure 4. Local affordance analysis of an example receptacle. LLM analyzes the contextual details from the scene graph and visual image to determine the potential functionality of the receptacle, assigning a more specific category.
@@ 3.1. Context-induced Affordance Analysis
To simplify the problem setting, we assume that a vanilla scene graph $S$ has already been constructed from an RGBD scanning sequence for the given scene, where each node represents an object instance in the scene and two nodes will be connected with an edge if they are close enough. Moreover, several attributes of the object instance are stored in each node, including the instance ID, category label, located room label, and rearrangement type (carriable, receptacle, or others), where the rearrangement type follows the definitions from the Vocabulary Mobile Manipulation (OVMM) (Yenamandra et al., 2023) Challenge. In the task, only carriable objects can be rearranged, receptacle objects serve as targets for placing carriable objects, and other objects do not interact with the agent during the task but can provide contextual information for other objects. The relationship between two nearby objects is also categorized into three different types, including near, on, and support, and stored on the connecting edge. Note that if a scene graph is not given, it can be constructed using established methods, with detailed guidance available in the supplementary material.
To help with the inference, we further associate each node with a representative RGB frame selected from the input RGBD scanning sequence to provide a visual context. In more detail, for each object instance, we first calculate the number of its projected pixels across all RGB-D images, creating an object-to-image-pixel matrix ($N*M$), where $N$ is the number of the object instance and $M$ is the number of frames. To select the most informative frame, we identify all the neighboring objects and sum up their projected pixel numbers together with that of the current instance on each frame. The frame with the maximal number of projected pixels is selected as the key frame for the node, as it contains the most visual information of the surrounding context of the instance to be considered.
@@
Local context refers to the neighboring information stored in the scene graph $S$. For each receptacle and other object in the $S$, we input all the textual contents related to its neighborhood and structure it using a predefined template. This template, along with its associated keyframe, forms the basis of a prompt designed for a Large Language Model (LLM) to analyze the object’s functionality through Chain of Thought reasoning (Wei et al., 2022). Consequently, the output is systematically formatted to detail the object in its Geometry & Position, Relationships, Functionality, and Fine-grained Category.
Figure 4 shows the local affordance analysis result of an example receptacle. In this example, the LLM first analyzes the geometric features of a table based on the visual information provided by the corresponding image. It combines this information with contextual information given by the scene graph to infer that the table is positioned against the wall at the entrance of the living room. Then, leveraging the position information and the supporting objects of the table in the scene graph, the LLM infers that the table can be likely used for placing outdoor wear. In the functionality module, based on the preceding analysis, the LLM further deduces that the table is suitable for placing items such as shoes that are typically found near entrances, as well as keys and other items that are temporarily stored when entering or leaving. As a conclusion, the LLM assigns a new fine-grained category, ”Entryway organizing console”, to the table based on the above analysis.
@@
Although our vanilla scene graph $S$ can effectively describe the surrounding context of a receptacle, the functionality of the receptacle may also depend on objects that are far away thus without connecting by an edge in $S$. For example, a sofa and a TV have a strong functional relationship (people sit on the sofa to watch TV), but they are typically arranged at a considerable distance from each other. Therefore, to consider more global context in the scene, we first build a hierarchical structure of $S$ and use it to bridge objects that are spatially distant yet functionally interrelated. In this way, new semantic edges are added, and the information stored in each object node is also updated, resulting in a new affordance-enhanced graph (AEG), as shown in Figure 3. The details of the hierarchy construction process can be found in the supplementary material. Note that objects are first grouped into an intermediate granularity level called “area”, and then further merged into a room. Using this ”object-area-room” hierarchy to aggregate contextual information is more efficient than aggregating directly from objects to rooms. This method allows for layered, structured storage of aggregated information, reserving important details.
@@
With the “object-area-room” hierarchy, we first use LLM to aggregate contextual information for each area separately. For each area, we collect the node information stored in $S_{l}$ of all the objects inside the area and design a prompt for LLM to generate the following content based on a fixed template: Description and Fine-grained name. Therefore, each area will be associated with the above information together with the list of the objects inside this area. We then integrate all the information from the areas within a room and use it as the global context for the room to extract distant context for all the receptacles and fixed objects within the room.
@@
For each receptacle, we input the basic spatial information template description, its local affordance analysis as well as the global context of the room to the LLM. With these inputs and a designed prompt, LLM can extract the distant but meaningful context from the global information and add these contexts as semantic edges to the SG. These context outputs will be in a format of a list of objects, indicating objects within the room that are functionally related to the given receptacle, as well as the specific functional relationship between these objects and the given receptacle. After obtaining this meaningful context, we add semantic directed edges from the given receptacle to related objects, reflecting their functional relations.
Figure 5. An example of placement scoring with the LLM. We set a fixed standard in the prompt for the LLM to refer to when rating: 100 for perfect placement, 0 for wrong placement, and 50 for placements difficult to judge.
Affordance updation.
@@
In the end, we use LLM to update context-induced affordance based on the previous analysis, the additional global context, as well as the semantic edges, with the same output format as in the local context affordance. In this way, our context-induced affordance includes both surrounding spatial context and distant semantic context, resulting in a more comprehensive understanding of receptacle’s functionality in this room. We refer to the final scene graph with updated node information and newly added semantic edges as Affordance Enhanced Graph (AEG) and use it for further misplacemnet detection as well as rearrangement planning.